OpenStudy (anonymous):

1) Equilibruim price: You can sell 280 products per week if they are marked at $15 each, but only 240 each week if they are marked at $18 each. Your supplier is prepared to sell you 320 products each week if they are marked at $20 each and 220 each week if they are marked $12 each. Find the demand equation, supply equation and equilibrium price for this product, using variables P for sales piece and Q for sales quantity. Now I NEED TO KNOW HOW TO DO IT AND I KNOW THE ANSWERS Demand equation: q=-13.33p+480 Supply equation: 12.5p+70 Equilibrium Price: $15.87 I know how to get -13.33p but...

OpenStudy (anonymous):

formula for demand is q=mp+b and I cant remember supply's formula :$ I need to know how to get B

OpenStudy (anonymous):

And to get m i used x2-x1/y2-y1

OpenStudy (anonymous):

treat both the supply and demand equation like linear equations of the form: y=mx+b where quantity is y, and price is x

OpenStudy (anonymous):

so both have the same linear equation? right?

OpenStudy (anonymous):

but how do i get b?

OpenStudy (anonymous):

once you have the slope substitute in a quantity and its price to find the constant aka y -intercept

OpenStudy (anonymous):

I remember i took a price from the both demand prices and then i used the formula but i cant seem to pull it off :/

OpenStudy (anonymous):

then set them equal to each other and solve for the price

OpenStudy (anonymous):

uhm... I dont get it...srry but can you make B?

OpenStudy (anonymous):

use the point when the price is 15 dollars and the quantity demand is 280 280=15*-13.33+B B=280+200 B=480

OpenStudy (anonymous):

oh god thnx by any chance do you know the equilibrium price formula?

OpenStudy (anonymous):

set the demand equation equal to the supply equation -13.33p+480=12.5p+70, solve for p

OpenStudy (anonymous):

ok thnx you are a life saver, i gave you a medal ;D

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

I may ask later another question in like 10 min if you have the time